OpenStudy (tomektomeknyc):

How to solve this: sin theta x cot^2 theta x sec^2 theta Answers: a) sec theta b) csc theta c) tan theta sec theta d) cos theta

OpenStudy (amoodarya):

\[\sin \theta \times \cot^2 \times \sec^2 \theta =\\sin \theta \times \frac{ \cos^2 \theta }{ \sin^2 \theta } \times \frac{ 1 }{ \cos^2 \theta } =\]

OpenStudy (tomektomeknyc):

thank you